# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ Abstract
This paper presents a novel approach to four fundamental problems in the field of cosmological perturbation theory. Firstly, the issue of gauge dependence has been addressed by demonstrating the existence of unique and gauge-invariant quantities corresponding to the actual perturbations. Secondly, the formation of primordial structures after decoupling of matter and radiation is dependent on the existence of local fluid flows resulting from local pressure gradients. To take pressure gradients into account, it is necessary to consider both the energy density and the particle number density. Thirdly, the novel relativistic perturbation theory applies to an open, flat, and closed Friedmann-Lemaître-Robertson-Walker universe. The derivation of the novel perturbation theory definitively reveals the inherent limitations of Newtonian gravitation as a framework for investigating cosmological density perturbations. Finally, the application of the perturbation theory to a flat universe demonstrates that, prior to decoupling, perturbations in both Cold Dark Matter and ordinary matter are coupled to perturbations in radiation. Therefore, the universe’s earliest structures formed only after decoupling, at which point local nonadiabatic random pressure fluctuations became a significant factor. Negative nonadiabatic pressure fluctuations resulted in a brief, rapid growth of density fluctuations until the total pressure fluctuations became positive. In contrast, positive nonadiabatic pressure fluctuations led to the formation of voids. Perturbations with masses about $2.2\times 10^{4}\,\text{M}_{\odot}$ became nonlinear already $60$ million years after the Big Bang, and perturbations with masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$ became nonlinear within about $600$ million years.
@@ I Introduction
The linearized Einstein equations and conservation laws of Friedmann-Lemaître-Robertson-Walker (flrw) universes are crucial to cosmology. They can explain the formation of all kinds of structures in the expanding universe, including stars, galaxies, and microwave background fluctuations.
Lifshitz [1] and Lifshitz & Khalatnikov [2] were the first researchers to develop a cosmological perturbation theory to study the evolution of density perturbations in the universe. They encountered the problem that the solutions of the linearized Einstein equations and conservation laws can be modified by linear coordinate transformations, which are known as gauge transformations. Consequently, the solutions are contingent upon the selection of a coordinate system, rendering them gauge-dependent and thus devoid of physical significance. This is the notorious gauge problem of cosmology, which demands urgent resolution. A multitude of potential solutions to this problem have been proposed in the literature. Stewart [3] noted in his paper, […] perturbation theory within general relativity is generally regarded as a somewhat arcane subject […] . This observation is further substantiated by the lack of progress observed to date in this field. The purpose of this paper is to elucidate the hitherto puzzling aspects of cosmological perturbation theory.
Before delving into an examination of the approaches outlined in the extant literature, it is imperative to acknowledge the fundamental criteria that a comprehensive and precise cosmological perturbation theory must fulfill. Primarily, as Bardeen elucidated in his seminal paper [4], a quantity signifying the real physical energy density perturbation must be manifestly gauge-invariant, that is, it must be defined independently of any reference system. Secondly, in the nonrelativistic limit, the linearized energy density constraint equation, when considered together with the linearized momentum constraint equation, must yield the time-independent Poisson equation of Newtonian theory with the aforementioned gauge-invariant quantity as the source term. Thirdly, in the nonrelativistic limit, it is essential that the freedom to select a coordinate system is present. That is to say, relativistic gauge transformations must converge to Newtonian gauge transformations in the nonrelativistic limit. Finally, and equally important, the perturbation theory must encompass all previously established and well-known physical solutions for large-scale density perturbations.
The present paper proposes a novel cosmological perturbation theory that has been demonstrated to satisfy all four of the aforementioned criteria. A review of the existing literature reveals that there is no theory that meets all four of the above criteria.
Bardeen [4] was the first to demonstrate that the linearized Einstein equations and conservation laws can be recast into evolution equations that have solutions independent of the choice of a coordinate system by using gauge-invariant quantities. In his paper, Bardeen constructed two distinct gauge-invariant quantities for the perturbation to the energy density, designated as $\epsilon_{m}$ and $\epsilon_{g}$. In the limit of small scales, these quantities converge to the usual gauge-dependent quantity, denoted as $\varepsilon_{{\scriptscriptstyle(1)}}$. Bardeen presupposed that on small scales, Newtonian gravitation is a valid description of cosmological density perturbations, and therefore, the gauge-dependent quantity, $\varepsilon_{{\scriptscriptstyle(1)}}$, becomes equal to its Newtonian counterpart. It is, however, a well-known fact that in Newtonian gravitation the coordinates of time and space can be chosen at will. Correspondingly, this must also be true in the nonrelativistic limit of a cosmological perturbation theory. It can thus be concluded that $\varepsilon_{{\scriptscriptstyle(1)}}$, which depends on the choice of coordinates in general relativity, is also dependent on the choice of coordinates in Newtonian gravitation. Thus, quantities that lack physical significance in general relativity are similarly devoid of physical meaning in Newtonian gravitation. It follows that the small-scale limit is not equivalent to the nonrelativistic limit. In consequence, the quantities $\epsilon_{m}$ and $\epsilon_{g}$ do not correspond to the actual density perturbations.
Bardeen’s paper has inspired a number of works, including those in Refs. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These researchers proposed alternative perturbation theories using gauge-invariant quantities that differ from those used by Bardeen.
The authors of Refs. [5, 6, 7, 9, 11] define gauge-invariant density perturbations by gradients of the energy density. For a pressureless fluid, their equations are identical to the standard perturbation equation derived from Newton’s theory of gravitation. This equation yields the well-known solutions of the standard equation (150) with $w=0$ and $\beta=0$, including the gauge mode. In addition to the fact that the solution is gauge-dependent, even in the nonrelativistic limit, the evolution of density perturbations in a pressureless fluid is precluded by the absence of pressure gradients. This is a consequence of the linearized momentum conservation laws. It can thus be concluded that the standard equation (150) with $w=0$ and $\beta=0$ cannot be taken to represent the nonrelativistic limit of a cosmological perturbation theory. Moreover, as indicated in Ref. [6], the constraint equations have not been integrated into the analysis. It is of crucial importance to acknowledge that the constraint equations represent relations between initial values, thereby playing a pivotal role in the analysis. In the absence of the constraint equations, the solutions to the system of dynamical equations and conservation laws lack physical significance. Furthermore, it is not possible to achieve a correct Newtonian limit without the constraint equations (for further details, see Sec. IX). In this paper, the complete set of Einstein equations and conservation laws is employed to develop a novel perturbation theory.
In order to interpret the gauge-invariant definition of the density perturbation $\overline{\delta\varepsilon}$, as proposed by Mukhanov et al. [8, 10], it is necessary to set the Hubble function $\cal{H}$ to zero in Eq. (7.47) in Ref. [10]. Nevertheless, this yields a Poisson-like equation that differs from the Newtonian Poisson equation, as the potential is time-dependent in the Poisson-like equation, as evidenced by equations (7.48) and (7.49) in Ref. [10]. Thus, $\overline{\delta\varepsilon}$ does not represent the actual physical density perturbation. Furthermore, the Hubble function cannot be set to zero, as this would violate the Friedmann equation (10a), which is inconsistent with the assumption that the universe is not empty. The present paper introduces a novel perturbation theory that achieves the time-independent Poisson equation and thus the correct nonrelativistic limit by considering zero pressure alone. This obviates the need to set the Hubble function to zero.
As has been stated in Refs. [12, 13], the linearized Einstein equations and conservation laws for density perturbations yield the usual evolution equation (150) derived from Newtonian gravitation when the pressure is zero. The current study will demonstrate that in a pressureless fluid, the linearized Einstein equations and conservation laws are reduced to a single equation: the time-independent Poisson equation of Newtonian gravitation.
The authors of Refs. [14, 15] present a method for the construction of gauge-invariant quantities in the synchronous reference system. However, the definition of gauge invariance presented in these works does not align with the conventional understanding of what constitutes a gauge-invariant quantity. As defined in Refs. [4, 3], a gauge-invariant quantity is independent of the choice of a coordinate system. Consequently, the definition of gauge-invariant quantities is formulated in this paper without the utilization of a coordinate system.
In their study of the evolution of density perturbations in a flat flrw universe in its radiation-dominated phase, Ma & Bertschinger [16] employ two distinct coordinate systems: the synchronous ($g_{0i}=0$) gauge and the conformal Newtonian gauge. The latter gauge is, in fact, a restricted synchronous gauge. Given the existence of gauge transformations between different synchronous reference systems, the solutions are dependent on the selected synchronous coordinates. Consequently, gauge modes emerge in the solutions, rendering them devoid of physical significance even when physical initial values are imposed. Their findings pertaining to the density fluctuation, represented by $\delta$, deviate between the synchronous and conformal Newtonian coordinates. In addition, Eq. (29) in Ref. [16] is flawed in two respects. Firstly, the factor $(1-3w)$ in the expansion term should, in fact, be $(2-3w)$, which is critical in the context of a radiation-dominated universe, where $w=\tfrac{1}{3}$. Secondly, it is crucial to replace the expression $k^{2}$ with $k^{2}/a^{2}$. This substitution is necessary because the background three-space metric of a flat flrw universe is $a^{2}\delta_{ij}$ rather than $\delta_{ij}$. This can be demonstrated by considering the line element (1) in Ref. [16], which illustrates that three-space is expanding. Once these corrections are made, the resulting equation is equal to the covariant divergence of Eq. (43) in this paper. The latter equation is correct, as demonstrated in Appendices B and C.
The aforementioned treatments resulted in a plethora of additional contributions to this field. See, e.g., Refs. [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In a review, Ellis [28] discusses the work of Lifshitz and Khalatnikov and other research on the subject. A review of the literature reveals a lack of consensus concerning the accuracy of the quantities employed to represent physical density perturbations.
This paper addresses the issue of gauge dependence in the context of cosmology. It demonstrates the existence of unique gauge-invariant quantities that correspond to the true physical density perturbations. Subsequently, it presents equations for the study of the evolution of density perturbations in universes with an open, flat, or closed flrw geometry. In light of the general covariance of general relativity, the evolution equations for gauge-invariant quantities are mathematically identical in all coordinate systems. Consequently, they yield the same solutions in any given coordinate system.
The novel perturbation theory, when applied to a flat flrw universe, reveals that prior to the decoupling of radiation and matter, Cold Dark Matter (cdm) perturbations were coupled to the radiation energy density perturbations in such a way that they could not clump together. Consequently, cdm proved to be ineffective in facilitating the process of structure formation subsequent to the decoupling of radiation and matter. The formation of the first stars, designated as Population iii stars [29], can be explained by the rapid growth of small density perturbations over a brief period of time. This rapid growth arises from local nonadiabatic random pressure fluctuations in the early universe following the decoupling. The James Webb Space Telescope (jwst), which was launched on December 25, 2021, is designed primarily for near-infrared astronomy and may therefore provide insights regarding the existence and nature of the earliest stars. The data from the jwst [30] suggest that the galaxies observed in the early universe appear to have grown to a size greater than anticipated, forming massive and compact structures at a rate exceeding the expectations of the hierarchical $\Lambda\textsc{cdm}$ structure formation paradigm. This may indicate that the existence of cdm is not supported by the data.
II A Novel Cosmological Perturbation Theory
This paper presents a novel cosmological perturbation theory, one that differs substantially from the perturbation theories in the extant literature, including the treatment of the cosmic fluid. The aim of this section is to provide an overview of the novel approach and to elucidate the method by which the new theory is constructed.
II.1 Cosmic Fluid
It is well established that in the era following the decoupling of matter and radiation, the pressure in the cosmological fluid is negligible in comparison to the energy density. Therefore, it can be concluded that the pressure can be disregarded in the Einstein equations and conservation laws of the background, i.e., the unperturbed, flrw universe. In the context of cosmology, it is commonly assumed that in the perturbed universe, the pressure and, therefore, local pressure gradients can also be neglected. However, it is well known that in the absence of local pressure gradients, local fluid flows cannot occur. This is equally true in the case of the cosmic fluid. As will be demonstrated in Sec. IX, the linearized momentum conservation laws indicate that, in the absence of pressure, local fluid flows will not occur and density perturbations will be static. It is thus crucial to consider pressure gradients in order to examine the evolution of density perturbations. Given that, subsequent to decoupling, the cosmic fluid is to be considered a perfect gas where the pressure is dependent on the mean kinetic energy of its particles, it is essential that the kinetic energy density be incorporated into the analysis. Both the pressure and the kinetic energy density are dependent on the particle number density and temperature. Consequently, in addition to the energy density, it is necessary to include the particle number density and the temperature in the equation of state. See Ref. [31], Sec. 2.10 on relativistic hydrodynamics, and Ref. [32]. As indicated in Ref. [33], the formation of structures in the universe requires a focus on the clumping of a set of particles rather than on the evolution of energy density instabilities. The approach presented in this paper yields novel insights into the evolution of density perturbations in both the pre-decoupling and post-decoupling eras. This is achieved by examining the evolution of energy density and particle number density perturbations, as well as their interactions. A detailed explanation of this can be found in Sec. XII.
II.2 Perturbation Theory
The equations governing the evolution of the unperturbed universe include three scalars: the energy density, the particle number density, and the expansion. As will be demonstrated in Sec. III, the perturbations to these three scalars transform under general linear coordinate transformations in the same way. This ultimately provides a solution to the gauge problem. As will be demonstrated in Sec. IV, there are only three sets of three gauge-invariant quantities that can be constructed from these scalars and their gauge-dependent perturbations. Each of these sets contains precisely one gauge-invariant quantity that is equal to zero. It will be shown that if the gauge-invariant perturbation of the expansion is zero, then it follows from the nonrelativistic limit in Sec. IX that the other two gauge-invariant quantities, denoted by $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, are the true physical perturbations of the energy density and the particle number density, respectively.
The subsequent step is to derive the evolution equations for $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$. In order to do this, it is necessary to employ a suitable coordinate system. In the context of general relativity, there is no a priori preferred coordinate system. Furthermore, the perturbations of the energy and particle number densities are both gauge-invariant, allowing for the freedom to choose any coordinate system. In order to facilitate the derivation of the aforementioned equations, it is typical to select a coordinate system that aligns with the specific characteristics of the problem in question. In Section Sec. V, a detailed justification is provided for the selection of synchronous coordinates. In Sec. VI, the Einstein equations, conservation laws, and their linearized counterparts with respect to this coordinate system are presented.
In Sec. VII, the synchronous reference system is employed to decompose the coupled system of linearized Einstein equations and conservation laws into three independent systems. These systems describe the evolution of tensor, vector, and scalar perturbations, respectively. In the extant literature, it is commonly assumed that the decomposition of the linearized equations can be achieved by decomposing only the perturbation of the metric tensor into its tensor, vector, and scalar parts. However, this is insufficient. Additionally, the perturbation of the spatial Ricci tensor, denoted by ${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}$, must similarly be decomposed into its tensor, vector, and scalar parts, as demonstrated in Ref. [34].
In the case of tensor and vector perturbations, the perturbation of the spatial Ricci scalar, or the spatial curvature perturbation, denoted by ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, is identically zero. In contrast, for scalar perturbations, the value of ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ is not zero. In addition, it will be demonstrated that the Newtonian potential is encapsulated within the expression for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$. This implies that only scalar perturbations are associated with density perturbations, specifically the physical density perturbations, represented by $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
In Sec. VIII, the system of equations (37) is derived which describes the evolution of scalar perturbations. This system incorporates not only the usual conservation laws for energy density and momentum, but also a conservation law for the particle number density. Furthermore, it is found that this system comprises an algebraic energy density constraint equation (37a) that incorporates the perturbed Ricci scalar ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, in addition to an evolution equation (37b) for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$. The latter equation is the covariant divergence of the three linearized momentum constraint equations (12b). In the nonrelativistic limit (i.e., the limit in which the pressure vanishes), the algebraic energy density constraint equation (37a) and the evolution equation (37b) for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, when combined, yield the time-independent Poisson equation of Newtonian gravitation with source term $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(\bm{x})$, as detailed in Sec. IX. Consequently, the constraint equations (37a)–(37b) are of great importance in the evolution of density perturbations and in the derivation of the nonrelativistic limit. The constraint equations presented in this specific formulation are notably absent from the extant literature on cosmological perturbation theories. This is the precise reason why an exact and correct nonrelativistic limit for a cosmological perturbation theory has not yet been discovered. The absence of a clearly defined nonrelativistic limit makes it impossible to interpret gauge-invariant quantities in previous perturbation theories. The issue is further discussed in Sec. IX, which addresses the nonrelativistic limit.
In Sec. X, the primary result is presented, namely the evolution equations (60) for the density fluctuations $\delta_{\varepsilon}:=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\delta_{n}:=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/n_{{\scriptscriptstyle(0)}}$ in closed, flat, and open flrw universes. As the quantities $\delta_{\varepsilon}$ and $\delta_{n}$ are gauge-invariant, the evolution equations will have the same mathematical form in all coordinate systems and yield solutions that are independent of the choice of a coordinate system.
In Sec. XII, the recently developed perturbation theory is applied to a flat flrw universe in its phases preceding and following the decoupling of radiation and matter. In order to proceed with the analysis, it is necessary to derive gauge-invariant thermodynamic quantities. These quantities will be presented in Sec. XI. Finally, in Sec. XIII, a summary of the results is provided.
The paper is accompanied by three appendices. In Appendix A, the primary result, Eqs. (60), is derived using the background equations (10), and the equations for scalar perturbations (37). The latter system represents a novel result and is indispensable for the proposed methodology. A proof of its correctness is provided in Appendices B and C. In Appendix C, the systems of equations (10) and (37) are employed once more. In this instance, however, the evolution equations (154) and (156) are recovered for the gauge-dependent quantity defined by $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, in a flat flrw universe. From these equations, the well-known physical solutions for large-scale perturbations in a flat flrw universe are derived. These include the sets $\{t,t^{1/2}\}$ for the radiation-dominated universe and $\{t^{2/3},t^{-5/3}\}$ for the universe following the decoupling of matter and radiation. A comparison of Eqs. (85) and (155a) and Eqs. (109) and (157a) reveals that the novel perturbation theory encompasses all previously established and well-known physical solutions for large-scale density perturbations. This constitutes a significant result, as the aforementioned solutions are widely documented in the literature. It follows that a comprehensive and accurate perturbation theory must also yield these solutions.
The homogeneous part of Eqs. (154a) and (156a) constitutes the conventional relativistic equation (150). Evidently, the latter equation is incomplete and has the gauge mode $\propto t^{-1}$ as solution. As will be demonstrated in Sec. IX, in the nonrelativistic limit, relativistic gauge transformations converge to Newtonian gauge transformations. Consequently, the quantity $\delta$ is also dependent on the choice of coordinates in Newtonian gravitation. Accordingly, it possesses no physical meaning. It can be concluded that the conventional relativistic evolution equation (150), is inadequate for examining the evolution of density fluctuations in the universe.
In the extant literature, the relativistic equation (150) is derived from Newtonian gravitation under two conditions: first, the expansion of the universe must be taken into account, and second, the equation of state must be nonrelativistic. The presence of the gauge mode in the solution of the aforementioned equation, when derived from the Newtonian theory of gravity, is the reason why Newtonian gravitation is not applicable to the study of density fluctuations in an evolving universe.
III Gauge Problem of Cosmology

### v3

# Earliest Structures in the Universe can be explained by a Relativistic Cosmological Perturbation Theory

@@ Abstract
A relativistic cosmological perturbation theory for the Friedmann-Lemaître-Robertson-Walker universe is presented that explains the masses and formation times of the first structures in our universe. First, it is shown that, without a coordinate system being used, quantities intended to represent energy density and particle number density perturbations can be defined in only one way. The Newtonian limit, where the pressure becomes zero, proves that these quantities are indeed the perturbations of the energy density and the particle number density. Then, after selecting a reference frame, a perturbation theory will be formulated based on these quantities. This formulation considers the local perturbation to the spatial curvature resulting from a density perturbation, the local fluid velocity due to pressure gradients caused by the self-gravity of the density perturbation, and entropy perturbations, all of which are necessary for structure formation. Pressure perturbations consist of two components: an adiabatic component caused by the density perturbation itself, and a random, nonadiabatic component resulting from the rapid, chaotic transition to the era when matter and radiation were decoupled. Immediately after decoupling, negative nonadiabatic pressure perturbations in various density perturbations enabled their rapid growth over a short period of time. This brief period ended when the total pressure perturbation became positive. Subsequently, the density perturbations gradually grew toward their nonlinear phase, which was reached early on.
@@ I Introduction
Structures in our universe, such as stars, galaxies, and black holes, form through the gravitational growth of small density perturbations. Therefore, it is reasonable to expect that a cosmological perturbation theory based on general relativity — the most accurate theory of gravity to date — would explain the masses and formation times of the first structures. There are two main reasons why such a theory does not yet exist. First, the gauge problem of cosmology has not yet been solved. Second, particle number density has not been considered, so the pressure gradients necessary for fluid flow — which is required for structure formation after matter and radiation decoupled — cannot be taken into account. This article resolves both issues.
The study of the evolution of density perturbations by way of linearization of the Einstein equations and conservation laws with respect to a flat Friedmann-Lemaître-Robertson-Walker (flrw) universe was pioneered by Lifshitz [1] and Lifshitz & Khalatnikov [2]. They observed that solutions to the linearized Einstein equations and conservation laws are dependent on the choice of a coordinate system, i.e., the solutions are gauge-dependent. Consequently, these solutions can be changed by linear coordinate transformations, also known as gauge transformations. This means that these solutions have no physical meaning, which is a problem known as the gauge problem of cosmology. Solving this long-standing problem is the only way to obtain physically significant solutions from the linearized Einstein equations and conservation laws. Doing so is paramount to explaining how the earliest structures [3, 4] in the universe formed from small density perturbations.
Physical phenomena such as density, temperature, pressure, and entropy perturbations do not depend on the choice of a coordinate system. Accordingly, the corresponding mathematical quantities describing these phenomena must be defined without using a reference frame. In other words, these quantities must be gauge-invariant.
Two different pioneering methods to address the gauge problem of cosmology have been proposed in the literature. In his seminal article, Bardeen [5] introduced gauge-invariant quantities intended to represent energy density perturbations as linear combinations of gauge-dependent quantities. Building on Bardeen’s work, Mukhanov et al.[6, 7] made significant contributions to the subject. Ellis et al.[8, 9, 10] present a second method that takes advantage of the fact that the gradient of a gauge-dependent quantity is gauge-invariant. For a comprehensive overview of this approach and its further extension, see Tsagas et al.[11].
Unfortunately, these approaches have some disadvantages. Firstly, they allow for different definitions of gauge-invariant quantities. Since these approaches lack a Newtonian limit, it is impossible to determine which gauge-invariant quantity corresponds to an energy density perturbation. Secondly, these approaches do not account for the particle number density. Thirdly, metric perturbations are explicitly present in the evolution equations for density perturbations. Finally, these approaches have not been extended to open and closed flrw universes.
Many new ideas have been added to this field. See Ellis’s $2017$ review article [12] and other publications, e.g., Refs. [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. These researchers proposed alternative perturbation theories using different gauge-invariant quantities that differ from those used by Bardeen and Mukhanov et al. However, the most fundamental question regarding which gauge-invariant quantity accurately and unambiguously represents a density perturbation remains unanswered.
This article presents a relativistic cosmological perturbation theory that does not have the aforementioned disadvantages. Firstly, it demonstrates that a gauge-invariant quantity describing a density perturbation can be defined in precisely one way. Secondly, in addition to energy density, particle number density [32] is considered. Both are crucial for structure formation. Finally, it is shown that perturbations in the spatial metric and its derivatives are included in the perturbations of the spatial Ricci scalar and expansion scalar. The use of these quantities significantly simplifies the system of linearized Einstein equations and conservation laws, leading not only to a perturbation theory that encompasses closed, flat, and open flrw universes but also allowing the Newtonian limit to be derived.
Unlike the approaches in Refs. [33, 34, 35, 36, 37, 38], this article does not introduce any ad hoc modifications to the theory of relativity, free parameters, Cold Dark Matter (cdm), or the peculiar velocities [39, 40, 41] of matter. However, the linearized Einstein equations and conservation laws naturally encompass the velocity of matter caused by pressure gradients, as well as the perturbed Ricci scalar caused by a density perturbation. These must be taken into account.
II Development and Overview of the Perturbation Theory
Sec. III explains why, in addition to energy density, particle number density must also be considered.
Sec. IV explains the gauge problem in cosmology and provides a solution by demonstrating that a gauge-invariant quantity that describes energy density perturbations can be defined in only one way. The same applies to the gauge-invariant perturbation of the particle number density. These definitions imply that the gauge-invariant perturbation to the expansion of the universe is equal to zero. In Sec. X, the Newtonian limit will be used to demonstrate that these quantities represent density perturbations.
To formulate a perturbation theory based on these quantities, a coordinate system is necessary. The reasons why a synchronous reference frame is the best choice for this task are explained in Sec. V.
In Sec. VI, the Einstein equations and conservation laws for flrw universes, as well as their linearized counterparts, are presented with respect to synchronous coordinates. Given that the evolution of density perturbations is considerably influenced by pressure perturbations, the covariant divergence of the fluid velocity, and the perturbation of the Ricci scalar, expressions are derived for these quantities.
Since we are working in synchronous coordinates, we can apply the decomposition theorem for symmetric, second-rank spatial tensors, such as the perturbed metric tensor and the perturbed spatial Ricci tensor, as well as the Helmholtz decomposition for the fluid velocity vector. These theorems are discussed in Sec. VII. According to these theorems, the linearized Einstein equations and conservation laws decompose into three independent systems of equations: one for tensor perturbations, one for vector perturbations, and one for scalar perturbations. In the literature, the decomposition theorem is only applied to the spatial metric tensor perturbation. However, this is insufficient. Since the perturbed Ricci scalar plays an important role in the evolution of density perturbations and in the Newtonian limit, the decomposition theorem must also be applied to the perturbed Ricci tensor.
Sec. VIII uses these theorems to show that only scalar perturbations are associated with density perturbations. In this case, the perturbed Ricci scalar contains the spatial derivatives of two independent potentials, while the perturbed expansion scalar contains their time derivatives. In the Newtonian limit, one of the potentials becomes the Newtonian potential while the other becomes irrelevant.
Since we are interested in density perturbations, we derive in Sec. IX a system of equations describing the evolution of scalar perturbations. This system contains three conservation laws: one for the energy density perturbation, one for the particle number density perturbation, and one for the momentum. It also contains two constraint equations. The algebraic energy density constraint equation contains the perturbed Ricci scalar, and the momentum constraint equation describes how the perturbed Ricci scalar evolves.
The Newtonian limit is reached when the velocities of the particles become negligible compared to the speed of light — that is, when the pressure approaches zero. Using this limit it is shown in Sec. X that the newly defined gauge-invariant quantities in Sec. IV represent perturbations of the energy and particle number densities. In other words, in the Newtonian limit, the momentum constraint equation ensures that the Newtonian potential and the energy density perturbation are independent of time. It also makes sure that the energy density constraint equation is identical to the Poisson equation of Newtonian gravity. Additionally, the well-known special relativistic equation relating rest mass density to energy density is obtained.
The Newtonian limit can only be obtained using the reformulated evolution equations for scalar perturbations given in Sec. IX. Since these equations are not presented in this form in the literature, previous cosmological perturbation theories have no Newtonian limit. Without a clearly defined Newtonian limit, it is impossible to demonstrate which gauge-invariant quantity represents a density perturbation. This article resolves this issue.
The Newtonian limit demonstrates furthermore that relativistic gauge transformations become Newtonian gauge transformations. Consequently, the conventional evolution equation for relative perturbations in a flat flrw universe filled with a nonrelativistic fluid — whether derived from general relativity or Newtonian gravity adapted to the expansion of the universe — also has gauge-dependent solutions devoid of physical meaning. Consequently, Newtonian gravity is not suitable for studying density perturbations in the universe, regardless of their scale. This means that no distinction can be made between density perturbations inside and outside the horizon.
The perturbation theory is presented in Sec. XI. It describes the evolution of density perturbations in closed, flat, and open flrw universes. The theory incorporates the perturbed Ricci scalar, the pressure, the entropy, and the covariant divergence of the fluid velocity. This is demonstrated by the algorithm in Appendix A.
In Sec. XII, expressions are derived for the perturbations of entropy, temperature, and pressure to connect the perturbation theory to an equation of state for a cosmological fluid. It will be shown that pressure perturbations have two components: an adiabatic component caused by the density perturbation itself, and a nonadiabatic component.
In Sec. XIII, we apply the perturbation theory to a flat flrw universe in its radiation-dominated phase and after the decoupling of matter and radiation.
Finally, we summarize the results and draw conclusions in Sec. XIV.
The article is accompanied by three appendices. Appendix A presents an algorithm that, when executed with the assistance of a computer algebra program, derives the perturbation theory from the background Einstein equations and the system of equations for scalar perturbations. Since the latter system is a new result, its correctness is demonstrated in Appendices B and C. Finally, Appendix C compares the conventional perturbation equations with the perturbation theory developed in this article.
III Equations of State
We distinguish between two main eras of the universe: the period before the decoupling of matter and radiation and the post-decoupling era.
Before decoupling, the cosmological fluid can be treated as a relativistic fluid consisting of radiation, ordinary matter, and cdm, where the contributions of matter and cdm to the pressure are negligible and the radiation pressure is of the same order of magnitude as the energy density. See, e.g., Ref. [6], Eq. (5.49), Ref. [7], Eq. (1.78), and Ref. [13], § V-1.
The conditions in the universe became significantly different after decoupling of matter and radiation. This is because the pressure in the cosmological fluid has become negligible compared to its energy density. Therefore, pressure can be neglected in the Einstein equations and conservation laws describing the evolution of a flrw universe. However, pressure perturbations cannot be ignored when studying structure formation. The self-gravity of a local density perturbation creates pressure gradients. These gradients then cause particle flows, thereby increasing a local density perturbation. The linearized momentum conservation law shows that the absence of pressure perturbations implies an absence of fluid flows. This results in static density perturbations. Therefore, pressure perturbations are crucial to understanding how structures in the universe formed. After decoupling, the cosmological fluid can be treated as a perfect nonrelativistic gas. In this case, pressure depends on the average kinetic energy density of the particles. Therefore, it is essential to include, next to the rest energy density, also the kinetic energy density in the analysis. Both pressure and kinetic energy density depend on the particle number density and temperature. Because of this, particle number density and temperature must be included in the equations of state. See Ref. [42], Sec. 2.10 on relativistic hydrodynamics, and Refs. [32, 43]. This results in random, nonadiabatic density perturbations, which are crucial for the formation of structures in the universe.
The aforementioned considerations are consistent with thermodynamic principles, namely that energy density $\varepsilon$ and pressure $p$ are functions of particle number density $n$ and temperature $T$, i.e.,
$\varepsilon=\varepsilon(n,T),\quad p=p(n,T).$ (1)
$p=p(n,\varepsilon),$ (2)
which will be used in the subsequent analysis.
IV Gauge Problem and its Solution